OpenStudy (anonymous):

6 – x – x = 18 6 6 9 –9 WILL GIVE MEDAL FOR CORRECT ANSWER!!!!

OpenStudy (anonymous):

its : 6 -2x = 18 -2x = 18 -6 -2x = 12 x = -6